OpenStudy (anonymous):

I'm Adding Right? 4x + 14 = -14 x =

TheSmartOne (thesmartone):

no

TheSmartOne (thesmartone):

you have to subtract.

TheSmartOne (thesmartone):

what is -14-14=?

OpenStudy (ganpat):

4x + 14 = -14 4x = -14 - 14 4x = -28 x = (-28)/4

OpenStudy (anonymous):

You have to do the inverse.

OpenStudy (anonymous):

It shows +, but you have to subtract.

OpenStudy (anonymous):

So -14 - 14 = ?

OpenStudy (anonymous):

@pootypie

OpenStudy (anonymous):

0

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

Use a calculator.

OpenStudy (andrewthecookie):

0 lol

OpenStudy (anonymous):

Sorry -28

OpenStudy (anonymous):

Yes!!!

OpenStudy (anonymous):

Then you divide -28 \(\div\) 4

TheSmartOne (thesmartone):

^ exactly :P

OpenStudy (anonymous):

>:P

OpenStudy (andrewthecookie):

-7

OpenStudy (anonymous):

Which Would Be -7

OpenStudy (anonymous):

@AndrewTheCookie Please don't give direct answers,

OpenStudy (andrewthecookie):

>>:P

OpenStudy (anonymous):

Yes! That is your answer! :)

OpenStudy (andrewthecookie):

ANSWER!!! YOU GAVE THE ANSWER ALREADY!!! GOD!

OpenStudy (andrewthecookie):

caps*

OpenStudy (anonymous):

No I didn't...

OpenStudy (andrewthecookie):

i mean u explained it